# Knowledge Homophily in Large Language Models

### v1

@@ Abstract.
Large Language Models (LLMs) have been increasingly studied as neural knowledge bases for supporting knowledge-intensive applications such as question answering and fact checking. However, the structural organization of their knowledge remains unexplored. Inspired by cognitive neuroscience findings, such as semantic clustering and priming, where knowing one fact increases the likelihood of recalling related facts, we investigate an analogous knowledge homophily pattern in LLMs. To this end, we map LLM knowledge into a graph representation through knowledge checking at both the triplet and entity levels. After that, we analyze the knowledgeability relationship between an entity and its neighbors, discovering that LLMs tend to possess a similar level of knowledge about entities positioned closer in the graph. Motivated by this homophily principle, we propose a Graph Neural Network (GNN) regression model to estimate entity-level knowledgeability scores for triplets by leveraging their neighborhood scores. The predicted knowledgeability enables us to prioritize checking less well-known triplets, thereby maximizing knowledge coverage under the same labeling budget. This not only improves the efficiency of active labeling for fine-tuning to inject knowledge into LLMs but also enhances multi-hop path retrieval in reasoning-intensive question answering. Our code is available at https://github.com/utkarshxsahu/kgc.
Structured Knowledge, Large Language Model, Homophily
@@ 1. Introduction
Large Language Models (LLMs) have emerged as powerful neural knowledge bases by encoding vast amounts of world knowledge within their neural parameters (Kadavath et al., 2022; Pezeshkpour, 2023). This neural-embedded knowledge enables LLMs to produce contextually relevant and factually rich responses, supporting real-world applications such as fact checking (Lin et al., 2021) and question answering (Yang et al., 2018; Wang et al., 2024). To better explore this neural knowledge base, researchers have devised knowledge checking methods to investigate the knowledge patterns of LLMs (AlKhamissi et al., 2022; Zheng et al., 2023) and leveraged the derived insights to guide knowledge-intensive tasks, including adaptive retrieval (Yao et al., 2024; Zhang et al., 2024), knowledge editing (Shi et al., 2024), and hallucination detection (Si et al., 2023).
@@
Figure 1. LLM is prompted on individual triple facts , which are aggregated to obtain entity-level knowledgeability scores. When visualized, these scores reveal the knowledge homophily pattern, where topologically close entities form distinct high-knowledge (red) and low-knowledge (blue) communities. Note that graph layout is visualized by the ForceAtlas2 algorithm (Jacomy et al., 2014) to preserve topological proximity.
@@
Despite various knowledge patterns identified previously  (Kadavath et al., 2022; Pezeshkpour, 2023; Zheng et al., 2024), little attention has been given to whether LLMs’ knowledge exhibits structural organization. In fact, in cognitive neuroscience (Liu et al., 2025), several works have highlighted the semantic clustered patterns of the neural knowledge in human brain networks: (i) semantic clustering in memory recall, where people tend to retrieve related words together (e.g., recalling “dog, cat, horse” in sequence) (Manning and Kahana, 2012; Bousfield and Sedgewick, 1944), and (ii) homophily brain networks, where regions with similar functions or inputs are more likely to connect (Sporns, 2012). Analogously, we hypothesize that LLMs also exhibit a similar knowledge homophily pattern, i.e., they tend to possess similar levels of knowledge about conceptually related entities, as illustrated in Figure 1 by checking GPT-3.5’s knowledge about triplets from WD50K dataset. Discovering this phenomenon provides valuable insights into the structural knowledge organization in LLMs and the design of knowledge-intensive task solutions. For example, by estimating the knowledgeability of concepts based on related concepts, one can identify factually weaker areas, thereby enabling more efficient knowledge labeling for knowledge injection and retrieval as we conducted in Section 4.
@@
Given the existence of knowledge homophily in other disciplines and its potential implications, this paper uncovers this innovative pattern and designs graph-based machine learning models to exploit neighborhood information to predict knowledgeability scores and discover less-known regions. These insights then guide efficient labeling for LLM fine-tuning and improve knowledge retrieval for multi-hop reasoning in question answering. Our contributions are:
@@ 2. Related Work
Knowledge Checking for LLMs as Knowledge Bases (KBs). As model parameters keep scaling, LLMs have evolved from task-specific executors to general-purpose agents and neural knowledge bases capable of storing and providing factual information for knowledge-intensive applications (Petroni et al., 2019; Roberts et al., 2020). However, unlike traditional KBs with structured schemas (Vrandečić and Krötzsch, 2014), the knowledge stored in LLM is encoded implicitly and non-interpretably within its parameters. This inherent lack of transparency creates a critical need for methods to verify what these models “know” and to ensure their reliable utilization, motivating numerous work on knowledge checking. These methods can be broadly categorized by their goals: verifying factual accuracy (Lin et al., 2021; Hendrycks et al., 2020), assessing self-awareness (Kadavath et al., 2022; Tian et al., 2023), and gauging knowledge coverage and consistency of LLMs against their own internal knowledge and external knowledge in other formats (Luo et al., 2023; Mallen et al., 2022). While valuable, they primarily focus on the content perspective of LLMs’ internal knowledge instead of the structural perspective reviewed next.
Pattern Understanding of LLMs as Knowledge Bases. With LLMs increasingly used as neural knowledge bases, existing studies have sought to uncover the underlying structured knowledge patterns encoded in their parameters from two perspectives. The first group investigates where knowledge is stored within the neural parameters of LLMs. For example, studies have shown that feed-forward layers can act as key-value memories for specific facts (Geva et al., 2020) and that factual associations are often localized and editable within these mid-layer modules, sometimes referred to as “knowledge neurons” (Meng et al., 2022; Dai et al., 2021). The second focuses on how knowledge is structurally organized.  (Mruthyunjaya et al., 2023) evaluated LLMs on structural properties like symmetry, hierarchy, and path-following, showing that models fail on complex relational tests. These structural analyses of knowledge patterns suggest the existence of implicit structures within LLMs, yet the homophily property remains underexplored.
3. Knowledge Homophily Discovery
In this section, we compute triplet-level knowledgeability scores and aggregate these measures to obtain entity-level knowledgeability scores. After that, we assess homophily by quantifying the knowledgeability differences between neighboring entities in Section 3.2.1 and qualitatively visualizing their scores in Section 3.2.2, thereby revealing the knowledge homophily phenomenon.
3.1. Knowledgeability Computation
@@ 3.1.2. Calculating Entity Knowledgeability
Given the above calculated triplet knowledgeability, we obtain knowledgeability of the entity $v_{i}$ by aggregating from all triplets involving $v_{i}$ (Jia et al., 2019; Rings et al., 2022):
@@ 3.2. Homophily Computation and Analysis
Furthermore, we evaluate whether topologically close entities share similar knowledgeability, i.e., the homophily of entity knowledgeability $\mathcal{H}(v_{i})$. Inspired by existing homophily computation (Wang and Derr, 2021; Ma et al., 2021), we compute knowledgeability homophily as one minus the average absolute difference in knowledgeability between central node $v_{i}$ and its neighbors $\mathcal{N}(v_{j})$ in the knowledge graph:
(2) $\small\mathcal{H}(v_{i})=1-\frac{1}{|\mathcal{N}(v_{i})|}\sum_{v_{j}\in\mathcal{N}(v_{i})}|\mathcal{K}(v_{i})-\mathcal{K}(v_{j})|$
@@
where a smaller difference between neighboring entities, $|\mathcal{K}(v_{i})-\mathcal{K}(v_{j})|$, leads to a higher homophily value $\mathcal{H}(v_{i})$. We empirically quantify triplet and entity knowledgeability scores and analyze homophily patterns both quantitatively and qualitatively. We evaluate five representative LLMs: GPT-3.5, 4o, Gemini-2.5 Flash, LLaMA3.3-70B, and DeepSeek-V3. These models are assessed across five knowledge graphs: MVPKG (Mou et al., 2024), T-Rex (Elsahar et al., 2018), PharmKG (Zheng et al., 2021), WD50K (Galkin et al., 2020), and CoDEx-S (Safavi and Koutra, 2020). Among them, T-Rex, WD50K, and CoDEx-S represent general Wikipedia knowledge, whereas PharmKG8K and MVPKG focus on biomedical and political science. Our graph visualizations in Figure 1/3 employ the ForceAtlas2 (Jacomy et al., 2014) to position topologically close nodes visually close, enabling an intuitive assessment of whether they share similar knowledgeability scores.
@@
Figure 2. (a): Distribution of node knowledgeability homophily for each dataset; (b): Average knowledge homophily across datasets and LLMs with black line showing Citeseer benchmark (0.74) for high homophily in graph analysis.
3.2.1. Quantitative Analysis of Knowledge Homophily
Figure 2(a)/(b) presents the node homophily distribution and the average graph homophily across several knowledge graphs. In Figure 2(a), these distributions are all right-skewed and peak around 0.8, suggesting the majority portion of nodes and their neighbors tend to share similar knowledgeability scores. This high homophily property enhances graph machine learning in node-level prediction, such as node classification and regression (Zhu et al., 2021), and therefore inspires regression to predict entities’ knowledge scores in Section 4.1. Furthermore, incorporating temporal information into MVPKG results in a slight left shift. This slightly decreasing neighbor homophily score might indicate that the temporal dimension introduces greater complexity and finer distinctions in knowledgeability between the nodes and their neighbors. In addition, we compute the average graph homophily by averaging across all nodes in Figure 2(b). Across datasets and LLMs, the scores mostly exceed the homophily of the Citeseer (Song et al., 2022; Wang et al., 2022) benchmark (horizontal line), a classic homophily network for node classification. This consistency across diverse settings indicates that the observed homophily generally aligns with the conventional notion of a “high-homophily” network.
@@ 3.2.1. Quantitative Analysis of Knowledge Homophily
Furthermore, we compare the calculated knowledge homophily against a degree-matched random baseline. For each node $v$ instead of using its actual neighborhood $\mathcal{N}(v)$, we create a random “peer group” $\widehat{\mathcal{N}}(v)$ by randomly sampling the same number of entities $(\widehat{\mathcal{N}}(v)=|\mathcal{N}(v)|)$ from the entire graph. Then, a homophily score is calculated as the difference between the central node $v$ and this randomly sampled group. The true neighborhood homophily was found to be statistically significant than this random baseline (100 trials per dataset, two-tailed z-test, $p<0.01$), with extremely narrow 99% confidence intervals. The results in Figure 3(a) show that across datasets, the homophily calculated from true neighbors exceeds the random baseline, confirming that it is not a statistical artifact or random fluctuation, but rather a robust and intrinsic structural property of LLMs’ internal knowledge organization.
@@ 3.2.2. Qualitative Analysis of Knowledge Homophily
Figure 3(b) shows a sampled T-Rex subgraph with nodes colored by entity-level knowledgeability $\mathcal{K}(v)$. A geopolitical neighborhood (countries) forms a compact and uniformly high-knowledgeability region with small intra-neighborhood deviations, confirming the high knowledgeability homophily. A historical football neighborhood (event/players) is similarly coherent but at lower knowledgeability. Despite varying means, both exhibit knowledge homophily.
Figure 3. (a) Neighboring nodes possess similar knowledgeability scores than randomly sampled nodes, indicating that homophily is a intrinsic structural property of LLMs; (b) Entities with their distinct knowledgeability levels $\mathcal{K}(v)$ indicated by node color (Red = High, Blue = Low).
@@ 4. Knowledge Homophily Application
After identifying the knowledge homophily pattern in Section 3, where topologically proximate entities exhibit similar knowledgeability scores, we apply this finding to two knowledge-intensive tasks: (1) homophily-aware knowledge checking for more efficient fine-tuning, and (2) homophily-aware knowledge retrieval for enhanced question answering. The core idea is to train a GNN-based model to estimate entity-level knowledgeability based on their neighbor knowledgeability and thereby identify triplets in less-known regions. These triplets can then be prioritized either for knowledge checking, to maximize fine-tuning benefits when injecting knowledge into LLMs, or for retrieval, to complement missing knowledge of the downstream answer generator. Both tasks rely on knowledgeability estimation to pinpoint knowledge gaps.
@@ 4.1. Homophily-aware Knowledge Estimation
Given homophily is sufficient for high-utility GNN predictions (Ma et al., 2021), we design a GNN-based regression model to perform message-passing, aggregate neighboring entity embeddings, and predict previously unknown entity scores. Specifically, given a set of entities $\mathcal{V}^{\text{Train}}$ with known knowledgeability (by prompting LLMs), our goal is to train a GNN to estimate the knowledgeability of unseen entities. At each layer, the model performs Message Passing (MP) and Feature Transformation (TR), followed by regression:
@@
The initial node feature matrix is defined as $\widetilde{\mathcal{K}}^{0}=[\mathcal{X}_{1},\dots,\mathcal{X}_{v_{|\mathcal{V}|}}]^{\top}$, where each node feature $\mathcal{X}_{v_{i}}$ is a dense text embedding obtained from pretrained language models. After training on $\mathcal{V}^{\text{Train}}$, the model is further used to infer the knowledgeability scores for all entities in the knowledge graph, eliminating the need for resource/time-intensive knowledge probing via exhaustive LLM prompting. In the following, we will utilize these estimated knowledgeability scores of entities to either guide triplet selection for subsequent LLM fine-tuning in Section 4.2 or guide retrieval for answering reasoning-intensive multi-hop questions in Section 4.3. Due to space constraints, we only briefly describe the experimental setup and provide full details in Figure 4 in Appendix.

### v2

@@ Abstract.
Large Language Models (LLMs) have been increasingly studied as neural knowledge bases for supporting knowledge-intensive applications. However, the structural organization of their knowledge remains unexplored. Inspired by cognitive neuroscience, such as semantic clustering and priming, where knowing one fact increases the likelihood of recalling related facts, we investigate an analogous knowledge homophily pattern in LLMs. To this end, we map LLM knowledge into a graph representation through knowledge checking at triplet/entity levels. After that, we analyze the knowledgeability relationship between an entity and its neighbors, discovering that LLMs tend to possess a similar level of knowledge about relevant entities positioned closer in the graph. Motivated by this homophily principle, we propose a Graph Neural Network (GNN) regression model to estimate entity-level knowledgeability scores for triplets by leveraging their neighborhood scores. The predicted knowledgeability enables us to prioritize checking less well-known triplets, thereby maximizing knowledge coverage under the same labeling budget. This not only improves the efficiency of active labeling for fine-tuning to inject knowledge into LLMs but also enhances multi-hop path retrieval in reasoning-intensive question answering. Our code and supplementary is available at https://github.com/utkarshxsahu/kgc.
Large Language Model; Knowledge Checking; Homophily
journalyear: 2026copyright: ccconference: Proceedings of the Nineteenth ACM International Conference on Web Search and Data Mining; February 22–26, 2026; Boise, ID, USAbooktitle: Proceedings of the Nineteenth ACM International Conference on Web Search and Data Mining (WSDM ’26), February 22–26, 2026, Boise, ID, USAdoi: 10.1145/3773966.3779394isbn: 979-8-4007-2292-9/2026/02ccs: Computing methodologies Natural language processing00footnotetext: Equal contribution and co-first authors.
@@ 1. Introduction
Large Language Models (LLMs) have emerged as powerful neural knowledge bases by encoding vast amounts of world knowledge within their neural parameters (Kadavath et al., 2022; Pezeshkpour, 2023). This neural-embedded knowledge enables LLMs to produce contextually relevant and factually rich responses, supporting real-world applications such as fact checking (Lin et al., 2021) and question answering (Lei et al., 2025; Wang et al., 2024). To better explore this neural knowledge base, researchers have devised knowledge checking methods to investigate the knowledge patterns of LLMs (AlKhamissi et al., 2022; Zheng et al., 2023) and leveraged the derived insights to guide knowledge-intensive tasks, including adaptive retrieval (Yao et al., 2024; Zhang et al., 2024; Han et al., 2024), knowledge editing (Shi et al., 2024), and hallucination detection (Si et al., 2023).
@@
Figure 1. We check whether LLM knows about triple facts and aggregate them to obtain entity knowledgeability scores. The visualized entity-level scores reveal the knowledge homophily, where topologically close entities form distinct high/log-knowledge (red/blue) communities. Graph layout is by ForceAtlas2 (Jacomy et al., 2014) to preserve topological proximity.
@@
Despite various knowledge patterns identified previously  (Kadavath et al., 2022; Pezeshkpour, 2023; Zheng et al., 2024), little attention has been given to whether LLMs’ knowledge exhibits structural organization. In fact, in cognitive neuroscience (Liu et al., 2025), several works have highlighted the semantic clustered patterns of the neural knowledge in human brain networks: (i) semantic clustering in memory recall, where people tend to retrieve related words together (e.g., recalling “dog, cat, horse” in sequence) (Manning and Kahana, 2012; Bousfield and Sedgewick, 1944), and (ii) homophily brain networks, where regions with similar functions or inputs are more likely to connect (Sporns, 2012). Analogously, we hypothesize that LLMs also exhibit a similar knowledge homophily pattern, i.e., they tend to possess similar levels of knowledge about conceptually related entities, as illustrated in Figure 1 by checking GPT-3.5’s knowledge about triplets from WD50K dataset. Discovering this phenomenon sheds light on how knowledge in LLMs is structurally organized and informs solutions for knowledge-intensive tasks. In particular, estimating a concept’s knowledgeability from related concepts helps identify weaker regions, enabling more efficient labeling for knowledge injection and retrieval as shown in Section 4.
@@
Motivated by homophily in other disciplines (Wang and Derr, 2021; Ma et al., 2021), this paper uncovers this pattern in LLMs and develops graph models to predict knowledgeability. These predictions identify less-known regions to guide efficient fine-tuning and enhance retrieval for multi-hop question answering. Our contributions are:
@@ 2. Related Work
Knowledge Checking for LLMs as Knowledge Bases (KBs). LLMs have evolved into general-purpose agents and neural knowledge bases for knowledge-intensive applications (Petroni et al., 2019; Roberts et al., 2020). However, unlike prior knowledge bases with explicit schemas (Vrandečić and Krötzsch, 2014), LLM knowledge is implicitly encoded and largely non-interpretable. This lack of transparency motivates the need to verify what LLMs “know” and ensure their reliable use. Existing knowledge checking methods can be categorized into verifying factual accuracy (Lin et al., 2021; Hendrycks et al., 2020), assessing self-awareness (Kadavath et al., 2022; Tian et al., 2023), and evaluating knowledge coverage and consistency against internal or external sources (Luo et al., 2023; Mallen et al., 2022). While effective, they focus on knowledge content rather than structure.
Structured Understanding of LLMs as Knowledge Bases. Existing structured understandings of LLM knowledge focus on model parameters from two perspectives. The first examines where knowledge is stored, showing that feed-forward layers act as key–value memories for factual knowledge (Geva et al., 2020), with factual associations often localized and editable in mid-layer “knowledge neurons” (Meng et al., 2022; Dai et al., 2021). The second investigates how knowledge is structurally organized. (Mruthyunjaya et al., 2023) evaluates properties such as symmetry, hierarchy, and path-following, revealing failures in complex relational reasoning. Despite exposing implicit structure in LLM knowledge, the role of homophily remains largely unexplored.
@@
3. Knowledge Homophily Discovery
This section investigates knowledge homophily. We first compute triplet-level knowledgeability and aggregate it into entity-level scores, then assess homophily by measuring knowledgeability differences between neighboring entities in Section 3.2.1 and qualitatively visualizing these scores in Section 3.2.2.
3.1. Knowledgeability Computation
@@ 3.1.2. Calculating Entity Knowledgeability
Given the above triplet knowledgeability, we obtain $v_{i}$ entity knowledgeability by aggregating scores of all triplets involving $v_{i}$ (Rings et al., 2022):
@@ 3.2. Homophily Computation and Analysis
We evaluate whether topologically close entities share similar knowledgeability, i.e., the homophily of entity knowledgeability $\mathcal{H}(v_{i})$. Following (Ma et al., 2021), we compute knowledgeability homophily as one minus the average absolute knowledgeability difference between central node $v_{i}$ and its neighbors $\mathcal{N}(v_{j})$:
(2) $\small\mathcal{H}(v_{i})=1-\frac{1}{|\mathcal{N}(v_{i})|}\sum_{v_{j}\in\mathcal{N}(v_{i})}|\mathcal{K}(v_{i})-\mathcal{K}(v_{j})|\vskip-0.96873pt$
@@
where a smaller difference between neighboring entities, $|\mathcal{K}(v_{i})-\mathcal{K}(v_{j})|$, leads to a higher homophily value $\mathcal{H}(v_{i})$. We empirically quantify triplet/entity-level knowledgeability and analyze homophily patterns both quantitatively and qualitatively. We evaluate five representative LLMs, GPT-3.5, 4o, Gemini-2.5 Flash, LLaMA3.3-70B, and DeepSeek-V3, across five knowledge graphs: MVPKG (Mou et al., 2024), T-Rex (Elsahar et al., 2018), PharmKG (Zheng et al., 2021), WD50K (Galkin et al., 2020), and CoDEx-S (Safavi and Koutra, 2020). T-Rex, WD50K, and CoDEx-S capture general Wikipedia knowledge, while PharmKG8K and MVPKG focus on biomedical and political domains. Graph visualizations in Figures 1 and 3 use ForceAtlas2 (Jacomy et al., 2014) to position topologically close nodes visually close, enabling an intuitive assessment of whether they share similar knowledgeability scores.
@@
Figure 2. (a): Homophily distribution of node knowledgeability; (b): Average knowledge homophily across datasets/LLMs with black dashed line showing a classic high homophily Citeseer (0.74) dataset for node classification (Wang and Derr, 2021).
3.2.1. Quantitative Analysis of Node/Graph Knowledge Homophily
Figure 2(a)/(b) shows node/graph-level homophily across multiple knowledge graphs. In Figure 2(a), node homophily distributions are right-skewed and peak near 0.8, indicating that most entities share similar knowledgeability with their neighbors. This high homophily is known to benefit node-level prediction tasks (Zhu et al., 2021), motivating our use of regression for entity knowledge estimation (Section 4.1). Incorporating temporal information in MVPKG causes a slight left shift. In addition, Figure 2(b) reports average graph homophily, which consistently exceeds that of the Citeseer benchmark (Song et al., 2022; Wang et al., 2022) across datasets/LLMs, indicating that the observed homophily aligns with the conventional “high-homophily” level (Ma et al., 2021).
@@ 3.2.1. Quantitative Analysis of Node/Graph Knowledge Homophily
We further compare knowledge homophily to a degree-matched random baseline by replacing each node’s true neighborhood $\mathcal{N}(v)$ with a randomly sampled peer group $\widehat{\mathcal{N}}(v)$ of the same size from the graph, and computing homophily relative to this group. True-neighborhood homophily is significantly higher than the random baseline (100 trials per dataset, two-tailed z-test, $p<0.01$) with tight 99% confidence intervals. As shown in Figure 3(a), this confirms that knowledge homophily is not a random artifact but an intrinsic structural property of LLMs’ knowledge organization.
@@ 3.2.2. Qualitative Analysis of Knowledge Homophily
Figure 3(b) visualizes a T-Rex subgraph colored by knowledgeability $\mathcal{K}(v)$. A geopolitical neighborhood forms a high-knowledge region, while a historical football cluster is similarly coherent but with lower knowledgeability. Despite varying means, the small intra-neighborhood deviations in both groups confirm strong knowledge homophily.
Figure 3. (a) Neighboring nodes possess similar knowledgeability scores to randomly sampled nodes ; (b) Entities with their distinct knowledgeability levels $\mathcal{K}(v)$ indicated by node color (Red = High, Blue = Low).
@@ 4. Knowledge Homophily Application
After discovering the knowledge homophily, where topologically proximate entities exhibit similar knowledgeability, we apply this insight to two knowledge-intensive tasks: (1) homophily-aware knowledge checking for efficient fine-tuning, and (2) homophily-aware knowledge retrieval for enhanced question answering. We train a GNN-based model to estimate entity-level knowledgeability from neighborhood signals and identify triplets in low-knowledge regions. These triplets are then prioritized for fine-tuning to maximize knowledge injection or for retrieval to complement missing knowledge in answer generation. Both tasks rely on knowledgeability estimation to pinpoint knowledge gaps.
@@ 4.1. Homophily-aware Knowledge Estimation
Given that homophily is a sufficient condition for high-utility GNN predictions (Ma et al., 2021), we design a GNN-based regression model to perform message-passing, aggregate neighboring embeddings, and predict unknown entity scores. Specifically, given a set of entities $\mathcal{V}^{\text{Train}}$ with known knowledgeability (by prompting LLMs), we train a GNN to estimate the knowledgeability of unseen entities. At each layer, the model performs Message Passing (MP) and Feature Transformation (TR), followed by regression:
@@
The initial node feature matrix is $\widetilde{\mathcal{K}}^{0}=[\mathcal{X}_{1},\dots,\mathcal{X}_{v_{|\mathcal{V}|}}]^{\top}$, where each node feature $\mathcal{X}_{v_{i}}$ is a dense textual embedding from pretrained language models. After training on $\mathcal{V}^{\text{Train}}$, the model is further used to infer the knowledgeability scores for all entities in the knowledge graph, eliminating the need for resource/time-intensive knowledge probing via exhaustive LLM prompting. We utilize the estimated entity knowledgeability scores to guide triplet selection for LLM fine-tuning (Section 4.2) and to guide retrieval for reasoning-intensive multi-hop QA (Section 4.3). Due to space constraints, we summarize the setup, and full details are in Figure 4 of Appendix.